OpenStudy (anonymous):

Find the LCM of each pair by listing multiples. 10, 45

OpenStudy (anonymous):

\(10 = 2 \times 5\) So, \(2\) and \(5\) are multiples of \(10\). Can you find that for \(45\) ??

OpenStudy (anonymous):

45=9x5 SO, 9 and 5 are multiples of 45.

OpenStudy (anonymous):

Good try..

OpenStudy (anonymous):

45 = 9X5 but 9 is again : 3x3.

OpenStudy (anonymous):

\[45 = 3 \times 3 \times 5\]

OpenStudy (anonymous):

3 and 5 are the multiples of 45, okay?

OpenStudy (anonymous):

Wait..

OpenStudy (anonymous):

whats the answer.